OpenStudy (anonymous):

9+x/7 9*7 = z/7*7 9*7=x/7*7 9*7=x*7/7*1 9*7=x*7/7 left is 9*7 = ? I am stuck here? Can anyone assist?

OpenStudy (anonymous):

Is your original equation\[9+\frac{ x }{ 7 }=0\]?

OpenStudy (anonymous):

No. The equation is 9=x/7

OpenStudy (anonymous):

Then x=9*7 which is 63.

OpenStudy (anonymous):

Yes...I have the answer. But I have not simplified the right side?

OpenStudy (anonymous):

\[\frac{ 7 }{ 7 }x=1x=x\]

OpenStudy (anonymous):

Can you explain the answer so I can understand?

OpenStudy (anonymous):

What is confusing you?

OpenStudy (anonymous):

I have all of this done, however my homework assignment is stating that I need to complete the right side. So should I add the entire answer? you can see all of my work up to that point.

OpenStudy (anonymous):

It looks like you have everything you need done... You've solved the equation for x, which is simplest form in this case.

OpenStudy (anonymous):

Thank you for your help....I wanted to be sure I understood..

OpenStudy (anonymous):

It seems that I have forgotten something? I perform multiplication at this point?